The Czech Republic has a gdp of 2000 billion koruny. The exchange rate is 20 koruny per u.s. dollar. The Czech population is 20 million. Caluculate the per capital gdp of the Czech Republic in u.s. dollars


1
Expert's answer
2021-11-08T17:21:35-0500

Since the Czech Republic has a GDP of 2,000 billion koruny.


The exchange rate is 20 koruny per U.S. dollar.


The Czech population is 20 million.


The per capita GDP of the Czech Republic = "\frac{GDP}{population}"

"=\frac{2,000 billion koruny }{ 20 million.}"


"=\frac{2,000,000 million koruny} {20 million}"

"=100,000 koruny"


The per capita GDP of the Czech Republic in US dollars


"=\frac{100,000 koruny} {20}"


"=\$5,000"


The per capita GDP of the Czech Republic in US dollars $5,000 per head.